# A Lagrangian Approach to the Inhomogeneous Incompressible Euler Equation

### v1

@@ 4 Lagrangian Analyticity of Solution
Now, thanks to (58), assuming for sufficiently small $t>0$, we have the following series expansion of $\ell$:
@@
Notice here $\ell^{(0)}=X(0,a)-a=0$, $\ell^{(1)}=u_{0}(a)$. Now straightforward computation yields:
@@
$\frac{d^{m}}{dt^{m}}\bigg\rvert_{t=0}\det\big(\delta_{ij}+\sum_{m=1}^{\infty}\partial_{j}\ell_{i}^{(m)}t^{m}\big)=0$
Using standard matrix identity, we have
@@
Here we remark that the expansion stops at $m=d$, eventually leading to a polynomial of $\nabla\ell^{(r)}$ with degree $d$.
@@
$P(0,0)=0,\quad\partial_{w}P(0,0)=-1\neq 0$
The implicit function theorem implies that locally around $(0,0)$, there exists an analytic function $\zeta_{*}(t)$ such that $P(\zeta_{*}(t),t)=0$. By uniqueness of solution of $P(t,\cdot)$, we conclude the formal series solution must coincide with analytic function $\zeta_{*}(t)$, which concludes the absolute convergence of series $\tilde{\zeta}$ in $t\in(0,T(\alpha^{*}))$, hence also the convergence of $\bar{\zeta}(t)$. Hence, Proof of the Lagrangian analyticity of IIE with convergence radius $t\in(0,T(\alpha^{*}))$ is complete. ∎

### v2

@@ 4 Lagrangian Analyticity of Solution
Now, assuming for sufficiently small $t>0$, we have the following formal series expansion of $\ell$:
@@
Notice here $\ell^{(0)}=X(0,a)-a=0$, $\ell^{(1)}=u_{0}(a)$. Straightforward computation yields:
@@
Here the expansion eventually leads to a polynomial of lower order $\nabla\ell^{(r)}$s’ enries with degree at most $m$.
@@
$P(0,0)=0,\quad\partial_{\zeta}P(0,0)=-1\neq 0$
The implicit function theorem implies that locally around $(0,0)$, there exists an analytic function $\zeta_{*}(t)$ such that $P(t,\zeta_{*}(t))=0$. By uniqueness of solution of $P(t,\cdot)$, we conclude the formal series solution must coincide with analytic function $\zeta_{*}(t)$, which concludes the absolute convergence of series $\tilde{\zeta}$ in $t\in(0,T(\alpha^{*}))$, hence also the convergence of $\bar{\zeta}(t)$. Hence, Proof of the Lagrangian analyticity of IIE with convergence radius $t\in(0,T(\alpha^{*}))$ is complete. ∎